if the sum of four consecutive integers is 86 what is the least of these integers
Answer by oberobic(2304)

Any 4 consecutive integers can be defined as w, x, y, and z.
But that makes for a really hard problem to solve.
A simpler way to define them is:
w = w
x = w+1
y = w+2
z = w+3
.
Now we have all 4 values in terms of 'w'.
.
We know the sum is 86.
.
w + (w+1) + (w+2) + (w+3) = 86
collect terms
4w + 6 = 86
subtract 6 from both sides
4w = 80
divide both sides by 4
w = 20
.
checking
20+21+22+23 = ??
86!
.
Answer:
The least of these numbers is 20.
.
Done.